# Differentially Private Zeroth-Order Methods for Scalable Large Language Model Fine-tuning

### v4

@@ A.1. Complete Proof
(43) $\begin{split}f_{\beta_{s}}(\hat{\boldsymbol{\theta}}_{T_{s}})&-f_{\beta_{s}}(\boldsymbol{\theta}^{*})\leq\left(\frac{1}{2\lambda}+\frac{1}{2\eta_{s}T_{s}}+\frac{1}{2T_{s}\lambda}\right)\left\|\boldsymbol{\theta}^{s-1}-\boldsymbol{\theta}^{*}\right\|^{2}\\ &+\eta_{s}\left(\frac{dc_{2}^{2}C^{2}PT\log(P/\delta)}{\epsilon^{2}n^{2}}+\frac{64d\beta_{s}^{2}L^{4}}{Pm}+\frac{\gamma^{2}}{Pm}+\frac{8dC^{2}}{ePm}\right).\\ \end{split}$
@@
(44) $\begin{split}f_{\beta_{s}}(\boldsymbol{\theta}^{s})-&f_{\beta_{s}}(\boldsymbol{\theta}^{*})\leq\left(\frac{1}{4\lambda\mu}+\frac{1}{4\eta_{s}T_{s}\mu}+\frac{1}{4T_{s}\lambda\mu}\right)\left(f_{\beta_{s-1}}(\boldsymbol{\theta}^{s-1})-f_{\beta_{s-1}}(\boldsymbol{\theta}^{*})\right)\\ &+\eta_{s}\left(\frac{dc_{2}^{2}C^{2}PT\log(P/\delta)}{\epsilon^{2}n^{2}}+\frac{64d\beta_{s}^{2}L^{4}}{Pm}+\frac{\gamma^{2}}{Pm}+\frac{8dC^{2}}{ePm}\right)\\ \leq&\left(\frac{1}{4\lambda\mu}+\frac{1}{4\eta_{s}T_{s}\mu}+\frac{1}{4T_{s}\lambda\mu}\right)\alpha_{s-1}\\ &+\eta_{s}\left(\frac{dc_{2}^{2}C^{2}PT\log(P/\delta)}{\epsilon^{2}n^{2}}+\frac{64d\beta_{S}^{2}L^{4}}{Pm}+\frac{\gamma^{2}}{Pm}+\frac{8dC^{2}}{ePm}\right).\\ \end{split}$
By setting $\eta_{s}=\alpha_{s}\cdot\min\left\{\frac{Pm}{7\gamma^{2}},\frac{Pme}{56dC^{2}},\frac{\epsilon^{2}n^{2}}{7dc_{2}^{2}C^{2}PT\log(P/\delta)},\frac{Pm}{448d\beta_{S}^{2}L^{4}}\right\}$ and $\lambda=\frac{7}{2\mu},\eta_{s}T_{s}=\frac{7}{2\mu}$, we have
(45) $f_{\beta_{s}}(\boldsymbol{\theta}^{s})-f_{\beta_{s}}(\boldsymbol{\theta}^{*})\leq\alpha_{s}.$
@@
(46) $f_{\beta_{S}}(\boldsymbol{\theta}^{S})-f_{\beta_{S}}(\boldsymbol{\theta}^{*})\leq\alpha_{S}.$
@@
(47) $\mathcal{O}\left(\left(\gamma^{2}+dC^{2}+d\beta_{S}^{2}L^{4}+\frac{dC^{2}P^{2}m\log(P/\delta)}{\epsilon^{2}n^{2}}\right)\cdot\frac{1}{\mu\alpha}\right).$
Lemma A.4.
@@
(48) $\begin{split}\mathbb{E}[||\nabla_{\mathbb{V}}F_{\beta}(\boldsymbol{\theta}_{t})&-\nabla_{\mathbb{V}}\hat{f}_{\beta}(\boldsymbol{\theta}_{t},\xi_{t+1})||^{2}]\\ \leq&\frac{dr_{s}c_{2}^{2}C^{2}PT\log(P/\delta)}{\epsilon^{2}n^{2}}+\frac{8dr_{s}C^{2}}{ePm}+\frac{64dr_{s}\beta^{2}L^{4}}{Pm}+\frac{\gamma^{2}}{Pm}.\end{split}$
Proof of Lemma A.4.
@@
Theorem A.5 (Restatation of Theorem 4.6).
@@
(49) $\phi_{s}(\boldsymbol{\theta}^{\prime S})-\phi_{s}(\boldsymbol{\theta}^{*})\leq\alpha.$
@@
(50) $\mathcal{O}\left(\left(\gamma^{2}+rdr_{S}C^{2}+rdr_{S}\beta_{S}^{2}L^{4}+\frac{rdr_{S}C^{2}P^{2}m\log(P/\delta)}{\epsilon^{2}n^{2}}\right)\cdot\frac{1}{\mu\alpha}\right).$
Proof of Theorem 4.6.
Similar in Theorem 4.4, we can obtain that
(51) $\begin{split}f_{\beta_{s}}(\boldsymbol{\theta}^{s})-&f_{\beta_{s}}(\boldsymbol{\theta}^{*})\leq\left(\frac{1}{4\lambda\mu}+\frac{1}{4\eta_{s}T_{s}\mu}+\frac{1}{4T_{s}\lambda\mu}\right)\left(f_{\beta_{s-1}}(\boldsymbol{\theta}^{s-1})-f_{\beta_{s-1}}(\boldsymbol{\theta}^{*})\right)\\ &+\eta_{s}\left(\frac{dr_{s}c_{2}^{2}C^{2}PT\log(P/\delta)}{\epsilon^{2}n^{2}}+\frac{64dr_{s}\beta_{s}^{2}L^{4}}{Pm}+\frac{8dr_{s}C^{2}}{ePm}+\frac{\gamma^{2}}{Pm}\right)\\ \leq&\left(\frac{1}{4\lambda\mu}+\frac{1}{4\eta_{s}T_{s}\mu}+\frac{1}{4T_{s}\lambda\mu}\right)\alpha_{s-1}\\ &+\eta_{s}\left(\frac{dr_{S}c_{2}^{2}C^{2}PT\log(P/\delta)}{\epsilon^{2}n^{2}}+\frac{64dr_{S}\beta_{S}^{2}L^{4}}{Pm}+\frac{8dr_{S}C^{2}}{ePm}+\frac{\gamma^{2}}{Pm}\right).\\ \end{split}$
By setting $\eta_{s}=\alpha_{s}\cdot\min\left\{\frac{Pm}{7\gamma^{2}},\frac{Pme}{56dr_{S}C^{2}},\frac{\epsilon^{2}n^{2}}{7dr_{S}c_{2}^{2}C^{2}PT\log(P/\delta)},\frac{Pm}{448dr_{S}\beta_{S}^{2}L^{4}}\right\}$ and $\lambda=\frac{7}{2\mu},\eta_{s}T_{s}=\frac{7}{2\mu}$, we have
(52) $f_{\beta_{s}}(\boldsymbol{\theta}^{s})-f_{\beta_{s}}(\boldsymbol{\theta}^{*})\leq\alpha_{s}.$
@@
(53) $f_{\beta_{S}}(\boldsymbol{\theta}^{S})-f_{\beta_{S}}(\boldsymbol{\theta}^{*})\leq\alpha_{S}.$
@@
(54) $\mathcal{O}\left(\left(\gamma^{2}+dr_{S}C^{2}+dr_{S}\beta_{S}^{2}L^{4}+\frac{dr_{S}C^{2}P^{2}m\log(P/\delta)}{\epsilon^{2}n^{2}}\right)\cdot\frac{1}{\mu\alpha}\right).$
A.2. More Experiment Results
In this section, we will provide the experiment settings and more experiment results that are not included in the main body of this paper.
Table 11. The hyperparameters grids used for RoBERTa-large experiments. All experiments use 6K steps for training.
Experiment Hyperparameters Values
DPZero Batch size 64
Learning rate $\left\{2e-5,1e-5,5e-6\right\}$
ZO scale parameter 1e-6
DP-ZOSO, DP-ZOPO Batch size 64
Learning rate $\left\{2e-4,1e-4,4e-5,2e-5,1e-5\right\}$
ZO scale parameter 1e-6 to 1e-5
Pruning rate ($r$) $\left\{0.5\%,1\%,2\%,4\%,8\%,100\%\right\}$
@@ A.2. More Experiment Results
Regulation parameter 5e-4
DP-SGD Batch size 64
Learning rate $\left\{1e-4,5e-4,1e-3,5e-3\right\}$
DP-prefix Batch size 64
Learning rate $\left\{1e-2,3e-2,5e-2\right\}$
@@
Table 12. The hyperparameters grids used for OPT experiments. All experiments use 6K steps for training.
Experiment Hyperparameters Values
ICL Examples 32
DPZero Batch size 16
Learning rate $\left\{1e-6\right\}$
ZO scale parameter 1e-4
DP-ZOSO, DP-ZOPO Batch size 16
Learning rate $\left\{4e-5,3e-5,2e-5,1e-5,5e-6\right\}$
ZO scale parameter 1e-4 to 4e-4
Pruning rate ($r$) $\left\{0.5\%,1\%,2\%,4\%,8\%\right\}$
Stage size 3
Regulation parameter 5e-4
Table 13. Experiments on RoBERTa-large (350M parameters). $r=0.5\%$ denotes the static pruning with pruning rate $r=0.5\%$ before fine-tuning and $r=0.5\%$* denotes the dynamic pruning with constant pruning rate $r=0.5\%$
@@
Table 14. Experiments of RoBERTa-large zeroth-order and first-order fine-tuned on SNLI with different pruning rates.
(a) Zeroth-order
Pruning Rate $\epsilon=2$ $\epsilon=4$ $\epsilon=8$ $\epsilon=\infty$
$r=0.5\%$ 70.3 78.6 81.4 80.5
$r=1\%$ 78.4 80.4 79.2 80.3
$r=2\%$ 78.2 80.4 78.9 79.6
$r=4\%$ 79.5 78.3 77.4 79.3
$r=8\%$ 77.3 78.6 79.0 78.6
$r=100\%$ 76.8 78.9 76.5 78.1
(b) First-order
Pruning Rate $\epsilon=2$ $\epsilon=4$ $\epsilon=8$ $\epsilon=\infty$
$r=0.5\%$ 85.0 85.8 87.0 86.8
$r=1\%$ 87.0 86.8 86.5 85.8
$r=2\%$ 86.6 86.6 86.7 88.2
$r=4\%$ 82.8 85.2 86.0 87.6
$r=8\%$ 84.6 86.1 86.7 86.9
$r=100\%$ 86.4 87.0 87.2 90.0
Table 15. Experiments on OPT-2.7B with privacy budget $\epsilon=4$. 1-2-4% $\ast$ denotes the incremental pruning strategy of with $r=$1-2-4%. Dynamic pruning outperforms static pruning in all five datasets.
@@
Table 16. Experiments on RoBERTa-large (350M parameters) with privacy budget $\epsilon=2$.
Task SST-2 SST-5 SNLI MNLI TREC
$r=0.5\%$ 88.5 46.5 70.3 70.1 83.4
$r=1\%$ 91.7 48.8 78.4 68.1 88.6
$r=2\%$ 88.4 47.1 78.2 71.4 89.6
$r=4\%$ 92.2 46.6 79.5 69.8 85.0
$r=8\%$ 92.5 46.6 77.3 69.3 85.4
$r=100\%$ 92.9 45.8 76.8 66.3 83.4
Table 17. Experiments on RoBERTa-large (350M parameters) with privacy budget $\epsilon=4$.
Task SST-2 SST-5 SNLI MNLI TREC
$r=0.5\%$ 92.2 50.5 78.6 71.4 81.0
$r=1\%$ 92.2 47.1 80.4 64.7 88.2
$r=2\%$ 91.6 47.7 80.4 71.5 87.0
$r=4\%$ 93.2 47.0 78.3 69.7 86.4
$r=8\%$ 92.5 48.4 78.6 69.7 84.4
$r=100\%$ 92.6 42.0 78.9 67.3 83.8
Table 18. Experiments on RoBERTa-large (350M parameters) with privacy budget $\epsilon=8$.
Task SST-2 SST-5 SNLI MNLI TREC
$r=0.5\%$ 92.2 49.8 81.4 71.6 85.4
$r=1\%$ 92.2 47.4 79.2 73.1 90.8
$r=2\%$ 91.8 49.4 78.9 68.6 85.4
$r=4\%$ 93.6 48.0 77.4 68.4 87.0
$r=8\%$ 92.4 48.0 79.0 69.7 84.4
$r=100\%$ 93.2 41.2 76.5 69.2 84.2

### v5

@@ A-A Complete Proof
$f_{\beta_{s}}(\hat{\boldsymbol{\theta}}_{T_{s}})$ $-f_{\beta_{s}}(\boldsymbol{\theta}^{*})\leq\left(\frac{1}{2\lambda}+\frac{1}{2\eta_{s}T_{s}}+\frac{1}{2T_{s}\lambda}\right)\left\|\boldsymbol{\theta}^{s-1}-\boldsymbol{\theta}^{*}\right\|^{2}$ (43)
$+\eta_{s}\left(\frac{dc_{2}^{2}C^{2}PT\log(P/\delta)}{\epsilon^{2}n^{2}}+\frac{64d\beta_{s}^{2}L^{4}}{Pm}+\frac{\gamma^{2}}{Pm}+\frac{8dC^{2}}{ePm}\right).$
@@
$f_{\beta_{s}}(\boldsymbol{\theta}^{s})-f_{\beta_{s}}(\boldsymbol{\theta}^{*})$ (44)
$\leq$ $\left(\frac{1}{4\lambda\mu}+\frac{1}{4\eta_{s}T_{s}\mu}+\frac{1}{4T_{s}\lambda\mu}\right)\left(f_{\beta_{s-1}}(\boldsymbol{\theta}^{s-1})-f_{\beta_{s-1}}(\boldsymbol{\theta}^{*})\right)$
$+\eta_{s}\left(\frac{dc_{2}^{2}C^{2}PT\log(P/\delta)}{\epsilon^{2}n^{2}}+\frac{64d\beta_{s}^{2}L^{4}}{Pm}+\frac{\gamma^{2}}{Pm}+\frac{8dC^{2}}{ePm}\right)$
$\leq$ $\left(\frac{1}{4\lambda\mu}+\frac{1}{4\eta_{s}T_{s}\mu}+\frac{1}{4T_{s}\lambda\mu}\right)\alpha_{s-1}$
$+\eta_{s}\left(\frac{dc_{2}^{2}C^{2}PT\log(P/\delta)}{\epsilon^{2}n^{2}}+\frac{64d\beta_{S}^{2}L^{4}}{Pm}+\frac{\gamma^{2}}{Pm}+\frac{8dC^{2}}{ePm}\right).$
By setting $\eta_{s}=\alpha_{s}\cdot\min\{\frac{Pm}{7\gamma^{2}}$, $\frac{Pme}{56dC^{2}}$, $\frac{\epsilon^{2}n^{2}}{7dc_{2}^{2}C^{2}PT\log(P/\delta)}$, $\frac{Pm}{448d\beta_{S}^{2}L^{4}}\}$ and $\lambda=\frac{7}{2\mu},\eta_{s}T_{s}=\frac{7}{2\mu}$, we have
$f_{\beta_{s}}(\boldsymbol{\theta}^{s})-f_{\beta_{s}}(\boldsymbol{\theta}^{*})\leq\alpha_{s}.$ (45)
@@
$f_{\beta_{S}}(\boldsymbol{\theta}^{S})-f_{\beta_{S}}(\boldsymbol{\theta}^{*})\leq\alpha_{S}.$ (46)
@@
$\mathcal{O}\left(\left(\gamma^{2}+dC^{2}+d\beta_{S}^{2}L^{4}+\frac{dC^{2}P^{2}m\log(P/\delta)}{\epsilon^{2}n^{2}}\right)\cdot\frac{1}{\mu\alpha}\right).$ (47)
Lemma 8.
@@
$\mathbb{E}[$ $||\nabla_{\mathbb{V}}F_{\beta}(\boldsymbol{\theta}_{t})-\nabla_{\mathbb{V}}\hat{f}_{\beta}(\boldsymbol{\theta}_{t},\xi_{t+1})||^{2}]$ (48)
$\leq$ $\frac{dr_{s}c_{2}^{2}C^{2}PT\log(P/\delta)}{\epsilon^{2}n^{2}}+\frac{8dr_{s}C^{2}}{ePm}+\frac{64dr_{s}\beta^{2}L^{4}}{Pm}+\frac{\gamma^{2}}{Pm}.$
Proof of Lemma 8.
@@
Theorem 6 (Restatation of Theorem 4).
@@
$\phi_{s}(\boldsymbol{\theta}^{\prime S})-\phi_{s}(\boldsymbol{\theta}^{*})\leq\alpha.$ (49)
@@
$\footnotesize\mathcal{O}\left(\left(\gamma^{2}+rdr_{S}C^{2}+rdr_{S}\beta_{S}^{2}L^{4}+\frac{rdr_{S}C^{2}P^{2}m\log(P/\delta)}{\epsilon^{2}n^{2}}\right)\cdot\frac{1}{\mu\alpha}\right).$ (50)
Proof of Theorem 4.
Similar in Theorem 2, we can obtain that
$f_{\beta_{s}}(\boldsymbol{\theta}^{s})-f_{\beta_{s}}(\boldsymbol{\theta}^{*})$ (51)
$\leq$ $\left(\frac{1}{4\lambda\mu}+\frac{1}{4\eta_{s}T_{s}\mu}+\frac{1}{4T_{s}\lambda\mu}\right)\alpha_{s-1}+\frac{64\eta_{s}dr_{S}\beta_{S}^{2}L^{4}}{Pm}$
$+\frac{\eta_{s}dr_{S}c_{2}^{2}C^{2}PT\log(P/\delta)}{\epsilon^{2}n^{2}}+\frac{8\eta_{s}dr_{S}C^{2}}{ePm}+\frac{\eta_{s}\gamma^{2}}{Pm}.$
By setting $\eta_{s}=\alpha_{s}\cdot\min\{\frac{Pm}{7\gamma^{2}}$, $\frac{Pme}{56dr_{S}C^{2}}$, $\frac{\epsilon^{2}n^{2}}{7dr_{S}c_{2}^{2}C^{2}PT\log(P/\delta)}$, $\frac{Pm}{448dr_{S}\beta_{S}^{2}L^{4}}\}$ and $\lambda=\frac{7}{2\mu},\eta_{s}T_{s}=\frac{7}{2\mu}$, we have
$f_{\beta_{s}}(\boldsymbol{\theta}^{s})-f_{\beta_{s}}(\boldsymbol{\theta}^{*})\leq\alpha_{s}.$ (52)
@@
$f_{\beta_{S}}(\boldsymbol{\theta}^{S})-f_{\beta_{S}}(\boldsymbol{\theta}^{*})\leq\alpha_{S}.$ (53)
@@
$\small\mathcal{O}\left(\left(\gamma^{2}+dr_{S}C^{2}+dr_{S}\beta_{S}^{2}L^{4}+\frac{dr_{S}C^{2}P^{2}m\log(P/\delta)}{\epsilon^{2}n^{2}}\right)\cdot\frac{1}{\mu\alpha}\right).$ (54)
A-B Hyperparameters
In this section, we will provide the experiment settings of RoBERTa-large, OPT-1.3b and OPT-2.7b. We use 6000 steps of training for classification and question-answering tasks. However, due to the complexity of text-generation tasks (translation and summarization), we use 11000 steps for training. In the following tables, BS denotes batch size, LR denotes learning rate, regularization denotes the regularization coefficient and $\beta$ denotes the ZO scale parameter.
TABLE X: The hyperparameters grids used for RoBERTa-large.
Method Hyper Values
FT BS 64
LR {1e-4, 5e-4, 1e-3, 5e-3}
FT-prefix BS 64
LR {1e-2, 3e-2, 5e-2}
prefix tokens 5
DPZero BS 64
LR {2e-5, 1e-5, 5e-6}
$\beta$ 1e-6
DP-ZOSO DP-ZOPO BS 64
LR {2e-4, 1e-4, 5e-6}
$\beta$ 1e-6 to 1e-5
pruning rate {0.5%,1%,2%,4%,8%,100%}
@@ A-B Hyperparameters
regularization 5e-4
TABLE XI: The hyperparameters grids used for OPT-1.3b.
Method Hyper Values
FT BS 8, 16
LR {1e-6, 5e-6, 1e-5, 5e-5, 1e-4}
FT-prefix BS 16
LR {1e-2, 3e-2, 5e-2}
@@
DPZero BS 16, 64
LR {6e-7, 8e-7, 1e-6, 2e-6}
$\beta$ 1e-6
DP-ZOSO DP-ZOPO BS 16, 64
LR {6e-7, 8e-7, 1e-6, 2e-6}
$\beta$ 1e-6 to 1e-5
pruning rate {0.5%,1%,2%,4%,8%,100%}
stage size 3, 5
regularization 5e-4
TABLE XII: The hyperparameters grids used for OPT-2.7b.
Method Hyper Values
FT BS 4, 8
LR {1e-6, 5e-6, 1e-5, 5e-5}
FT-prefix BS 4, 8
LR {1e-3, 5e-3, 1e-2}
prefix tokens 5
DPZero BS 16
LR {6e-7, 8e-7, 1e-6, 2e-6}
$\beta$ 1e-4
DP-ZOSO DP-ZOPO BS 16
LR {6e-7, 8e-7, 1e-6, 2e-6}
$\beta$ 1e-6 to 1e-5
pruning rate {0.5%,1%,2%,4%,8%,100%}
stage size 3, 5
regularization 5e-4
A-C More Experiment Results
In this section, we will provide more experiment results that are not included in the main body of this paper.
TABLE XIII: Experiments on RoBERTa-large (350M parameters) with privacy budget $\epsilon=2$.
Task SST-2 SST-5 SNLI MNLI TREC
$r=0.5\%$ 88.5 46.5 70.3 70.1 83.4
$r=1\%$ 91.7 48.8 78.4 68.1 88.6
$r=2\%$ 88.4 47.1 78.2 71.4 89.6
$r=4\%$ 92.2 46.6 79.5 69.8 85.0
$r=8\%$ 92.5 46.6 77.3 69.3 85.4
$r=100\%$ 92.9 45.8 76.8 66.3 83.4
TABLE XIV: Experiments on RoBERTa-large (350M parameters) with privacy budget $\epsilon=4$.
Task SST-2 SST-5 SNLI MNLI TREC
$r=0.5\%$ 92.2 50.5 78.6 71.4 81.0
$r=1\%$ 92.2 47.1 80.4 64.7 88.2
$r=2\%$ 91.6 47.7 80.4 71.5 87.0
$r=4\%$ 93.2 47.0 78.3 69.7 86.4
$r=8\%$ 92.5 48.4 78.6 69.7 84.4
$r=100\%$ 92.6 42.0 78.9 67.3 83.8
TABLE XV: Experiments on RoBERTa-large (350M parameters) with privacy budget $\epsilon=8$.
Task SST-2 SST-5 SNLI MNLI TREC
$r=0.5\%$ 92.2 49.8 81.4 71.6 85.4
$r=1\%$ 92.2 47.4 79.2 73.1 90.8
$r=2\%$ 91.8 49.4 78.9 68.6 85.4
$r=4\%$ 93.6 48.0 77.4 68.4 87.0
$r=8\%$ 92.4 48.0 79.0 69.7 84.4
$r=100\%$ 93.2 41.2 76.5 69.2 84.2
TABLE XVI: Experiments of RoBERTa-large zeroth-order and first-order fine-tuned on SNLI with different pruning rates.
(a) Zeroth-order
Pruning Rate $\epsilon=2$ $\epsilon=4$ $\epsilon=8$ $\epsilon=\infty$
$r=0.5\%$ 70.3 78.6 81.4 80.5
$r=1\%$ 78.4 80.4 79.2 80.3
$r=2\%$ 78.2 80.4 78.9 79.6
$r=4\%$ 79.5 78.3 77.4 79.3
$r=8\%$ 77.3 78.6 79.0 78.6
$r=100\%$ 76.8 78.9 76.5 78.1
(b) First-order
Pruning Rate $\epsilon=2$ $\epsilon=4$ $\epsilon=8$ $\epsilon=\infty$
$r=0.5\%$ 85.0 85.8 87.0 86.8
$r=1\%$ 87.0 86.8 86.5 85.8
$r=2\%$ 86.6 86.6 86.7 88.2
$r=4\%$ 82.8 85.2 86.0 87.6
$r=8\%$ 84.6 86.1 86.7 86.9
$r=100\%$ 86.4 87.0 87.2 90.0
Figure 11: The training loss on SNLI with RoBERTa-large of FT and DP-ZOPO with different privacy budgets ($\epsilon=2,4,8$) that are commonly considered in related works [1] [11] [13]. DP-ZOPO with larger privacy budget ($\epsilon=8$) has faster convergence rate. Under the same privacy budget, DP-ZOPO converges faster than FT in the early stage while FT can continue to converge more stably in the later stages.
TABLE XVII: Experiments on RoBERTa-large (350M parameters). Dynamic pruning with a constant pruning rate fails to improve optimization.
Task SST-2 SST-5 SNLI MNLI TREC
Small Privacy budget: $\epsilon=2$
static mask 92.5 48.8 79.5 71.4 89.6
Medium Privacy budget: $\epsilon=4$
static mask 93.2 50.5 80.4 71.5 88.2
Large Privacy budget: $\epsilon=8$
static mask 93.6 49.8 81.4 73.1 90.8
TABLE XVIII: Experiments on RoBERTa-large (350M parameters). $r=0.5\%$ denotes the static pruning with pruning rate $r=0.5\%$ before fine-tuning and $r=0.5\%$* denotes the dynamic pruning with constant pruning rate $r=0.5\%$.
@@ A-C More Experiment Results
TABLE XIX: Experiments on RoBERTAa-large. 1-2-4% $\ast$ and 1-2-4% denote the incremental pruning and dynamic pruning strategy separately, both with dynamic pruning rate $r$ increasing from 1% to 2% and 4%. Incremental pruning strategy outperforms dynamic pruning strategy in most dataset under all private settings.
Task SST-2 SST-5 SNLI MNLI TREC
Small Privacy budget: $\epsilon=2$
0.5-1-2% 89.4 48.7 79.2 71.3 86.6
0.5-1-2% $\ast$ 88.5 49.0 76.7 71.6 88.4
1-2-4% 92.8 48.2 77.6 71.3 85.2
1-2-4% $\ast$ 92.2 47.7 80.1 70.7 91.2
2-4-8% 89.6 46.4 75.3 72.0 91.4
2-4-8% $\ast$ 92.8 48.6 77.5 70.6 91.4
Medium Privacy budget: $\epsilon=4$
0.5-1-2% 93.0 48.0 79.2 71.3 86.6
0.5-1-2% $\ast$ 93.6 49.1 80.0 73.1 90.4
1-2-4% 92.8 48.2 79.4 71.2 87.4
1-2-4% $\ast$ 92.9 49.3 80.1 71.6 92.0
2-4-8% 92.0 47.9 78.7 74.9 88.6
2-4-8% $\ast$ 92.8 48.6 80.6 73.9 88.6
Large Privacy budget: $\epsilon=8$
0.5-1-2% 92.2 49.8 80.8 71.9 88.0
0.5-1-2% $\ast$ 93.6 48.5 79.2 74.8 89.2
1-2-4% 92.5 48.8 79.3 72.0 89.0
1-2-4% $\ast$ 92.9 48.1 79.4 72.4 88.0
2-4-8% 92.9 46.0 81.4 74.4 89.0
2-4-8% $\ast$ 92.8 48.6 81.6 74.8 91.8
TABLE XX: Experiments on OPT-2.7b with privacy budget $\epsilon=4$. 1-2-4% $\ast$ denotes the incremental pruning strategy of with $r=$1-2-4%. Dynamic pruning outperforms static pruning in all five datasets.
@@
TABLE XXI: OPT-1.3b and OPT-2.7b fine-tuned on SQuAD. DP-ZOPO has better QA accuracy than ICL, DPZero, DP-ZOSO, FT-prefix, even FT in all private settings.
(a) OPT-1.3b
Method $\epsilon=2$ $\epsilon=4$ $\epsilon=8$
Zero-shot 27.23 27.23 27.23
ICL 58.66 58.66 58.66
FT 71.71 72.24 75.30
FT-prefix 71.37 72.80 73.21
DPZero 48.38 52.17 47.46
DP-ZOSO 73.31 73.01 75.24
DP-ZOPO 73.71 76.00 75.24
(b) OPT-2.7b
Method $\epsilon=2$ $\epsilon=4$ $\epsilon=8$
Zero-shot 29.89 29.89 29.89
ICL 67.63 67.63 67.63
FT 77.35 77.41 78.35
FT-prefix 76.50 76.66 76.92
DPZero 60.18 63.46 65.12
DP-ZOSO 76.83 77.68 78.16
DP-ZOPO 78.83 78.03 79.03